A company that supplies ready-mix concrete receives, on average, six orders per
day.
(a) What is the probability that, on a given day:
(i) only one order will be received.
(ii) no more than three orders will be received.
(iii) at least three orders will be received.
(b) What is the probability that, on a given half-day, only one order will be
received?

Answer by Boreal(15235)

This is a Poisson distribution with number proportional to time, theoretically could be infinite and mass function.
P(1)=e^(-6)*6^1/1!=6e^(-6) or 0.0149
-
P(2)=e^(-6)*6^2/2!=0.0446
P(3)=0.08924
p(0)=e^(-6)=0.0025
so P(not greater than 3) is 0.1512 from the calculator rather than adding the four terms above
-
At least 3 means 1-P(0,1,2)=0.9380.
-
A half day is Poisson parameter 3, and the probability of receiving 1 is P(1)=e^(-3)*3^1/1!=3e(-3)=0.1494